OpenStudy (anonymous):

The percent P of the American population that graduated college between 1950 and 1990 is approximated by P=5.9556+1.492t+0.0056t² where t represent the calendar year with t=0 corresponding to 1950. When will the percent of college graduates exceed 25% of the population?

OpenStudy (turingtest):

0.25=5.9556+1.492t+0.0056t^2 0=5.7059+1.492t+0.0056t^2 the binomial theorem gives two answers for t in years. one will probably be negative and so can be thrown out. use the other value of t and add it to 1950 to get the correct year.

OpenStudy (anonymous):

binomial theorem ?